What measure of angle is swept by the minute hand during the duration of a class if the class starts at 2:16pm and ends at 2:47pm?
Answer by stanbon(75887)

What measure of angle is swept by the minute hand during the duration of a class if the class starts at 2:16pm and ends at 2:47pm?
---------
# of minutes: 2:47 - 2:16 = 31
--------------
# of degrees swept: (31/60)360 = 31*6 = 186 degrees
=========================
Cheers,
Stan H.